OpenStudy (anonymous):

What is the solution set of 7x^2 + 3x = 0? {0, 3/7} {0, -3/7} {0, -4/7}

OpenStudy (anonymous):

Take x common..

OpenStudy (anonymous):

how?

OpenStudy (campbell_st):

x (7x + 3) = 0 then x = 0 or -3/7

OpenStudy (anonymous):

x(7x + 3) = 0..

OpenStudy (anonymous):

Now when the product of two numbers is zero?? If one of the number must be zero.. So either x = 0 or (7x + 3) = 0 Find x from here...